how do you find the solution to -8-n+14=6
Answer by Alan3354(69443)

how do you find the solution to -8-n+14=6
------------------
I'll assume you want to solve for n.
-8-n+14=6
-n-8+14=6 Move things around
-n+6=6
-n = 0
n = 0